OpenStudy (anonymous):

Check to see if this equation: x^2+3x+2x+6 has rational roots using the Rational Root Theorem.

OpenStudy (anonymous):

we solved this already and saw that the roots were \(-2\) and \(-3\) which are integers, and are therefore rational

OpenStudy (anonymous):

The previous question wasn't the same. How did you get -2 and -3? @satellite73

OpenStudy (anonymous):

\[(x+2)(x+3)=0\] \[x+2=0\iff x=-2\] \[x+3=0\iff x=-3\]

OpenStudy (anonymous):

Oh I see what you did there okay